OpenStudy (anonymous):

write an equivalent expression without using grouping symbols. -(6x - 7)

OpenStudy (anonymous):

ARGH! I just wrote a whole explanation then I hit back and it disappeared. I'll do it quickly. Do you know what the distributive property is? 2(3x - 4) = 2*3x + 2*-4 = 6x - 8 ^--- distributive property being used Well, in this case, there's a hidden one in front of the parentheses. Anything multiplied by 1 is itself, so adding a one there doesn't change things. -(6x - 7) = -1(6x - 7) Distribute, -1*6x + -1*-7 = -6x + 7